You buy a computer directly from the manufacturer for R7332,00 and agree to repay it in equal instalments over three years at the end of each month form now . The interest rate is 10.7% per year, compounded monthly. How much interest will you pay in total?


1
Expert's answer
2021-07-29T17:01:27-0400


Given the following:

Future value (FV)= R7332

Rate (r)= 10.7% or 0.107

Time (n)= 3 years

The installmental parent is given by

"PV=\frac{FV}{[\frac{(1+r)^n-1}{r}]}"


"PV=\frac{7332}{[\frac{(1+0.107)^3-1}{0.107}]}"


"PV=R2200.18"




The picture above shows the payment schedule.

From the table to get interest for Year 2 "=" 2200.18"(" PV of year 1")" "\u00d7" 10.7%"=" 235.42

Balance to be paid"=" annual installment"+" deposit"+" interest

2200.18"+" 2200.18"+" 235.42"=" 4635.7

Interest for year 3"=" 4635.7"\u00d7" 10.7%"=" 496.02

Balance to be paid"=" 2200.18"+" 4635.7"+" 496.02"=" 7331.9


The total interest due in the payment is R731.44 "(" 235.42"+" 496.02")"